# Transonic accretion flow in the mini discs of a binary black hole system

### v1

@@ III Model equations
In this section, we aim to model the transonic accretion flow in the CPD within the stationary and axisymmetric BBH system (1). Hydrodynamics of the CPD are studied within the framework of general relativity [41]. We consider the fluid motion is restricted on the $z=0$ plane of the central source ($u^{z}=0$ and $\partial Q/\partial z=0$, where $Q$ represents any flow parameter). Additionally, we assume that the flow is steady ($\partial Q/\partial t=0$), axisymmetric ($\partial Q/\partial\phi=0$), and inviscid.
@@ III.1 Governing equations for CPD
The fundamental governing equations that describe the flow motion in the CPD are given by [42],
@@
where $v$ is the radial-component of the physical three-velocity in the co-rotating frame (CRF, i.e., a frame orbits with the same angular velocity $\Omega$ as the fluid) [43, 44], which takes negative values for accretion, and $\gamma_{v}=1/\sqrt{1-v^{2}}$ denotes the Lorentz-factor corresponding to $v$. Here, $\rho$ is the rest-mass density, $e$ ($=\rho+\Pi$, $\Pi$ is the internal energy density) is the total energy density, and $p$ is the isotropic pressure of the flow. The quantity $\Phi^{\rm eff}$ is called the effective potential of the accreting system, and it is obtained as,
@@
where $\lambda$ ($=-u_{\phi}/u_{t}$, $u_{\phi}$ and $u_{t}$ denote the $\phi$ and $t$ components of the covariant four-velocity $u_{k}$, respectively) is the specific angular momentum of the flow. Note that in the energy equation (6), we do not consider any radiative cooling processes due to the low mass accretion rate of the supermassive black holes [45, 46].
@@
where $E$ is known as the Bernoulli constant and $\mathcal{L}$ represents the bulk angular momentum per unit mass. Therefore, $\lambda$ ($=\mathcal{L}/E$) remains conserved along the flow direction. Using the normalization condition ($u^{k}u_{k}=-1$) of the four-velocity, we determine $u_{t}$ as,
@@
where $\Omega$ ($=u^{\phi}/u^{t}=\lambda V^{2}/x^{2}$) refers the angular velocity of the flow.
@@
where $H$ is the local half-thickness of the disc. The expression of $H$ is obtained by considering the hydrodynamic equilibrium along the transverse direction of the disc as [47, 48, 44],
@@
To solve the dynamical equations of the flow, we require an equation of state (EoS) that relates various thermodynamic quantities of the flow. We assume the fluid system consists of fully ionized plasma, which means fluid satisfies the charge neutrality condition $n_{p}=n_{e}\approx\rho/m_{p}$, where $n_{p}$ and $n_{e}$ represent the number densities of protons and electrons, respectively. Since for hot accretion flow (HAF), the plasma temperature can reach up to $\sim 10^{10-12}$ [46] near the horizon, we employ the relativistic EoS with variable adiabatic index $\Gamma$, as proposed in [49]. Accordingly, the internal energy density $e$ and isotropic pressure $p$ are defined as,
@@
where $m_{p}$ and $m_{e}$ are masses of proton and electron, respectively, and $\Theta$ [$=k_{B}T/(m_{e}c^{2})$, $k_{B}$ is the Boltzmann constant and $T$ is the flow temperature in Kelvin] is the dimensionless temperature. Here, the quantity $f$ is given by,
@@
where $\mathcal{K}$ is the entropy constant and $\chi=(f-1-m_{p}/m_{e})/(2\Theta)$. We determine the entropy accretion rate, which is also a constant of motion due to the adiabatic nature of the flow, and is given by [50, 51],
@@
where $N$ ($=\frac{1}{2}\frac{df}{d\Theta}$) refers the polytropic index.
@@
where $C_{s}$ ($=\frac{\Gamma p}{e+p}=\frac{2\Gamma\Theta}{f+2\Theta}$) is the adiabatic sound speed.
@@ III.2 Conditions for critical point
In case of transonic accretion flow, it must cross the event horizon ($x_{\rm H}$) with radial velocity close to the speed of light ($v\approx 1$), thereby satisfying the inner boundary condition at the event horizon [27, 28]. Usually, accreting matter begins its journey with negligible radial velocity ($v<<1$) at the outer edge of the accretion disc. As the flow moves inward, it must pass through at least one critical point ($x_{c}$) where it changes the sonic state from subsonic to supersonic. At $x_{c}$, the radial velocity gradient takes the indeterminate form $(dv/dx)_{x_{c}}=0/0$. Therefore, for the critical points, we find the following conditions:
@@ III.3 Radiative properties of accretion flow
We consider the relativistic thermal bremsstrahlung (free-free) emission from the CPD. For HAF, electron-electron emission dominates over electron-ion emission [52], and hence we use an approximate expression for the free-free emissivity, given by [53],
@@
where $Z$ represents the atomic number of the ion, which is taken as $1$ for the hydrogen plasma, $h$ is the Planck constant, and $\nu_{e}$ is the emission frequency. The term $\bar{g}_{\rm ff}$ denotes the thermally-averaged Gaunt factor, which does the quantum mechanical correction to the classical electrodynamics. Its value varies between $1$ to $5$, depending on the electron energy. However, following [45], we take $\bar{g}_{\rm ff}=1.2$ throughout this paper. Note that the second term in Eq. (22) includes both the relativistic corrections and electron-electron bremsstrahlung emission. Also, we assume a weak coupling between the electrons and ions, which leads to a temperature difference between them. There are many studies in the literature corresponding to the two-temperature accretion flows [54, 55, 56, 57, 58]. However, in accordance with [45], we use a simplified parameterized relation $T_{e}=T/10$.
The emission coefficient of accretion flow, as described in Eq. (22), is measured locally. However, for an observer at spatial infinity, this radiation must be red-shifted due to both the gravitational potential of the black hole and the Doppler effect resulting from the rotation of the disc. Here, we do not consider the light bending effect to the photons for simplicity. Also, the disc inclination angle with respect to the line of sight of the observer is taken as $45^{\circ}$ [59, 60, 52]. Under these assumptions, the observed frequency $\nu_{o}$ is obtained in terms of the emitted frequency $\nu_{e}$ as [61],
@@
[…]
where $x_{0}$ ($=x_{\rm H}=$ event horizon radius) and $x_{\rm edge}$ are the inner and outer edges of the CPD, respectively.
Bremsstrahlung radiation can only be observed when the emitting medium is optically thin; otherwise, photons will be destroyed through true absorption after undergoing multiple coherent scattering. The primary source of opacity is the Thomson scattering by the free electrons with opacity coefficient $\kappa_{s}=0.4~{}{\text{cm}^{2}~{}\text{gm}^{-1}}$ [62]. In a fully ionized medium, free-free absorption also significantly contributes to the photon opacity [62]. Using Eq. (22), we calculate the Rosseland mean opacity coefficient for the free-free absorption as [63],
@@
Here, we consider the typical length scale of the medium is the half-thickness ($H$) of the disc. Therefore, the effective optical depth of the disc is obtained by considering the aforementioned opacity sources as [64],
@@
where $\tau_{s}$ ($=\kappa_{s}\rho H$) is the scattering optical depth and $\tau_{a}$ ($=\kappa_{\rm R}^{\text{ff}}\rho H$) is the absorption optical depth.
@@ IV.1 Analysis of critical points
We already pointed out that the transonic accretion flow must pass through the critical points; therefore, we start our analysis by investigating how the behaviors of critical points depend on different global constants, such as specific angular momentum ($\lambda$), binary mass ratio ($q$), and binary separation ($z_{2}$). To do this, we first compute the flow temperature ($\Theta_{c}$) and radial velocity ($v_{c}$) at a critical point ($x_{c}$) for a given set of input parameters ($\lambda,q,z_{2}$) by solving the critical point conditions (21). Subsequently, using these results in Eq. (8), we calculate the specific energy ($E$) at $x_{c}$ for the same set of input parameters mentioned above. The obtained results are shown in Fig. 3, where the variation of $E$ as a function of $x_{c}$ is illustrated. In panel (a), we fix $(q,z_{2})=(1,500)$ and vary the angular momentum as $\lambda=3.3,3.35$, and $3.4$. In panel (b), we set $(\lambda,z_{2})=(3.5,500)$ and vary the mass ratio as $q=0.01,0.1$, and $1$. Finally, in panel (c), we choose $(\lambda,q)=(3.5,0.5)$ and vary the binary separation as $z_{2}=500,750$, and $1000$. The color codes corresponding to the different input parameters are indicated in each panel. In each curve, the saddle, nodal, and spiral-type critical points are represented by solid, dotted, and dashed lines, respectively. We observe that the flow energy remains constant in all cases when the critical points are located at large distances from the black hole. However, the flow energy changes significantly when the critical points are closer to the black hole. For each curve, the critical points are formed in a systematic order as saddle-nodal-spiral-nodal-saddle-spiral when the critical points move toward the outer edge of the disk. Moreover, we observe that the flow energy changes moderately with an increase in $\lambda$. However, it shows small variation with increases in $q$ and $z_{2}$. Interestingly, a broad range of energy values exists where the flow possesses two saddle-type critical points (as indicated by the horizontal dashed lines (magenta) in all the panels). It is important to note that such regions with multiple critical points are of particular interest in the study of shock-induced accretion solutions. For the shock scenarios [42, 65, 66, 60, 67, 68], a global solution can pass through both the inner (i.e., formed near the black hole) and outer (i.e., formed far from the black hole) critical points, provided that the relativistic shock conditions are satisfied. A detailed analysis of the shock solutions is presented in Section IV.3.
@@ IV.2 Effect of binary parameters on the transonic accretion solutions
In this section, we aim to investigate the effect of the secondary black hole (SBH) on the accretion solutions in the CPD. To determine the transonic accretion solutions, we calculate the critical point $x_{c}$ and its corresponding flow variables $v_{c}$ and $\Theta_{c}$ for a given set of input parameters ($q,z_{2},\lambda,E$) by using Eqs. (21) and (8). Next, using the calculated parameters ($\Theta_{c},v_{c}$) at $x_{c}$ as the initial boundary conditions, we first integrate Eqs. (16) and (18) from $x_{c}$ to the disc inner edge $x_{0}$, and then from $x_{c}$ to the disc outer edge $x_{\text{edge}}$. Finally, by combining both segments of the solutions, we obtain the resulting global accretion solution corresponding to the same set of input parameters ($q,z_{2},\lambda,E$). In this work, the inner and outer edges of the accretion disc are taken as $x_{0}=0.05$ and $x_{\text{edge}}=350$, respectively. It is essential to mention that there is no qualitative difference in the characteristics of the accretion solutions if the inner edge position is chosen more closer to the horizon ($x_{\rm H}=0$) instead of $x_{0}=0.05$. In Fig. 4, we depict the typical accretion solutions (i.e., variation of the Mach number ($M=v/C_{s}$) as a function of the radial coordinate ($x$)) for different values of $q$ and $z_{2}$. Here, we choose the input flow parameters as $\lambda=3.5$ and $E=1.01$. The left panels represent the accretion solutions corresponding to the mass ratio $q=0.01,0.1$, and $1$ with binary separation $z_{2}=500$. Similarly, the right panels denote the accretion solutions for $z_{2}=500,750$, and $1000$ with $q=0.5$. In all cases, the solutions that pass through the outer critical points ($x_{\text{out}}$) are globally extended from $x_{\text{edge}}$ to $x_{0}$. In contrast, the solutions that pass through the inner critical points ($x_{\text{in}}$) are truncated at certain radii ($x_{t}$) in between $x_{\text{in}}$ and $x_{\text{out}}$. Accretion solutions of this kind are usually referred as A-type solutions [42, 59, 60, 66, 52, 68]. Interestingly, for A-type solution topology, the solution that passes through $x_{\text{out}}$ may connect to the solution that passes through $x_{\text{in}}$ via a standing shock transition, as the inner solution branch has a higher entropy content than the outer one [69]. The shock-induced accretion solutions will be examined latter in Section IV.3. It is evident from the figure that the nature of the accretion solutions does not change with an increase in either of the binary parameters $q$ and $z_{2}$. The positions of the critical points associated with these accretion solutions are summarized in Table 1. This table shows that $x_{\text{in}}$ increases, while $x_{\text{out}}$ decreases as $q$ increases at a given $z_{2}$. On the other hand, when $z_{2}$ increases at a fixed $q$, $x_{\text{in}}$ and $x_{\text{out}}$ move oppositely compared to the previous case.

### v2

@@ III Model equations
In this section, we aim to model the transonic accretion flow in the CPD within the stationary and axisymmetric BBH system (1). Hydrodynamics of the CPD are studied within the framework of general relativity Rezzolla and Zanotti (2013). We consider the fluid motion is restricted on the $z=0$ plane of the central source ($u^{z}=0$ and $\partial Q/\partial z=0$, where $Q$ represents any flow parameter). It is important to note that the assumption of $u^{z}=0$ is a valid approximation when the binary mass ratio ($q$) is small and the binary separation ($z_{2}$) is large. This is because the gravitational force exerted by the secondary black hole on the accreting material surrounding the primary black hole is negligible compared to the force from the primary black hole. However, for moderate to large mass ratios (up to a maximum value of $1$), assuming $u^{z}=0$ becomes a rough approximation as the gravitational influence of the secondary black hole becomes effective. Nevertheless, to simplify our analysis, we adopt $u^{z}=0$ for all mass ratios. Furthermore, the metric exhibits time translation ($t$) and azimuthal ($\phi$) symmetries. Consequently, the BBH metric is associated with two Killing vectors along the $t$ and $\phi$ directions. Any matter living in this background must also adhere to these symmetries to maintain the stability of the entire system. This ensures the conservation of energy and the angular momentum flux associated with the flow. Under these circumstances, the matter near the primary black hole must have angular momentum aligned with the $Z$-axis, and hence matter will approximately lie on the $z=0$ plane as long as the mini disk is concerned. Additionally, we assume that the flow is steady ($\partial Q/\partial t=0$), axisymmetric ($\partial Q/\partial\phi=0$), and inviscid.
@@ III.1 Governing equations for CPD
The fundamental governing equations that describe the flow motion in the CPD are given by Dihingia et al. (2018a),
@@
where $v$ is the radial-component of the physical three-velocity in the co-rotating frame (CRF, i.e., a frame orbits with the same angular velocity $\Omega$ as the fluid) Lu (1985); Peitz and Appl (1997), which takes negative values for accretion, and $\gamma_{v}=1/\sqrt{1-v^{2}}$ denotes the Lorentz-factor corresponding to $v$. Here, $\rho$ is the rest-mass density, $e$ ($\rho+\Pi$, $\Pi$ is the internal energy density) is the total energy density, and $p$ is the isotropic pressure of the flow. The quantity $\Phi^{\rm eff}$ is called the effective potential of the accreting system, and it is obtained as,
@@
where $\lambda$ ($u_{\phi}/u_{t}$, $u_{\phi}$ and $u_{t}$ denote the $\phi$ and $t$ components of the covariant four-velocity $u_{k}$, respectively) is the specific angular momentum of the flow. Note that in the energy equation (8), we do not consider any radiative cooling processes due to the low mass accretion rate of the supermassive black holes Yarza et al. (2020); Yuan and Narayan (2014).
@@
where $E$ is known as the Bernoulli constant and $\mathcal{L}$ represents the bulk angular momentum per unit mass. Therefore, $\lambda$ ($\mathcal{L}/E$) remains conserved along the flow direction. Using the normalization condition ($u^{k}u_{k}=-1$) of the four-velocity, we determine $u_{t}$ as,
@@
where $\Omega$ ($u^{\phi}/u^{t}=\lambda V^{2}/x^{2}$) refers the angular velocity of the flow.
@@
where $H$ is the local half-thickness of the disc. The expression of $H$ is obtained by considering the hydrodynamic equilibrium along the transverse direction of the disc as Lasota (1994); Riffert and Herold (1995); Peitz and Appl (1997),
@@
To solve the dynamical equations of the flow, we require a relativistic equation of state (REoS) that relates various thermodynamic quantities of the flow (e.g., mass density, pressure, and temperature). We assume the fluid system consists of fully ionized plasma, which means fluid satisfies the charge neutrality condition $n_{p}=n_{e}\approx\rho/m_{p}$, where $n_{p}$ and $n_{e}$ represent the number densities of protons and electrons, respectively. Since for hot accretion flow (HAF), the plasma temperature can reach up to $\sim 10^{10-12}$K [^1] Note that the temperature at the inner edge of the disc exceeds the typical threshold for nuclear fusion in stellar cores, which lies in the range of $10^{7}-10^{8}$ K. However, the mass density of matter in the accretion disc is significantly lower than that in stellar cores. Moreover, the interaction time between nuclei is very short due to the short accretion timescale. For these reasons, despite the high temperature at the disc’s inner edge, the other physical conditions are not favorable for nuclear fusion to occur in the accretion disc. Yuan and Narayan (2014) near the horizon, we employ the relativistic EoS with variable adiabatic index $\Gamma$, as proposed in Chattopadhyay and Ryu (2009). Accordingly, the internal energy density $e$ and isotropic pressure $p$ are defined as,
@@
where $m_{p}$ and $m_{e}$ are masses of proton and electron, respectively, and $\Theta$ $k_{B}T/(m_{e}c^{2})$, $k_{B}$ is the Boltzmann constant and $T$ is the flow temperature in Kelvin] is the dimensionless temperature. Here, the quantity $f$ is given by,
@@
where $\mathcal{K}$ is the entropy constant and $\chi=(f-1-m_{p}/m_{e})/(2\Theta)$. We determine the entropy accretion rate, which is also a constant of motion due to the adiabatic nature of the flow, and is given by Chattopadhyay and Kumar (2016); Kumar and Chattopadhyay (2017),
@@
where $N$ ($\frac{1}{\Gamma-1}=\frac{1}{2}\frac{df}{d\Theta}$) refers the polytropic index. Note that, for simplicity, the ideal equation of state (IEoS) is often used, where $N$ (or $\Gamma$) is treated as a constant. However, in our case, it is a variable quantity, as the temperature exceeds $\sim 10^{10}$ K near the horizon Chattopadhyay and Ryu (2009); Dihingia et al. (2019). Using the above definition of $N$, one can obtain its profile provided the temperature profile is known, which can be found from Eq. (18).
@@
where $C_{s}$ ($\frac{\Gamma p}{e+p}=\frac{2\Gamma\Theta}{f+2\Theta}$) is the adiabatic sound speed.
The above flow equations are useful for determining the accretion solutions in the CPD. However, we mentioned earlier that in the absence of both the external field and the secondary black hole, the spacetime metric of a BBH system [Eq. (1)] reduces to that of a single Schwarzschild black hole of mass $m_{1}$ (see Eqs. (4) and (5)). Consequently, all the flow equations, namely Eqs. (9), (11), (12), (17), and (18), must reduce to the standard flow equations corresponding to an isolated Schwarzschild black hole. Therefore, one can expect that these equations provide accretion solutions similar to those reported in the literature for an isolated Schwarzschild black hole (Chattopadhyay and Ryu, 2009; Chattopadhyay and Kumar, 2016, and references therein).
@@ III.2 Conditions for critical point
In case of transonic accretion flow, it must cross the event horizon ($x_{\rm H}$) with radial velocity close to the speed of light ($v\approx 1$), thereby satisfying the inner boundary condition at the event horizon Liang and Thompson (1980); Abramowicz and Zurek (1981). Usually, accreting matter begins its journey with negligible radial velocity ($v<<1$) at the outer edge of the accretion disc. As the flow moves inward, it must pass through at least one critical point ($x_{c}$) where it changes the sonic state from subsonic to supersonic. At $x_{c}$, the radial velocity gradient takes the indeterminate form $(dv/dx)_{x_{c}}=0/0$. Therefore, for the critical points, we find the following conditions:
@@ III.3 Radiative properties of accretion flow
We consider the relativistic thermal bremsstrahlung (free-free) emission from the CPD. For HAF, electron-electron emission dominates over electron-ion emission Patra et al. (2024a), and hence we use an approximate expression for the free-free emissivity, given by Novikov and Thorne (1973),
@@
where $Z$ represents the atomic number of the ion, which is taken as $1$ for the hydrogen plasma, $h$ is the Planck constant, and $\nu_{e}$ is the emission frequency. The term $\bar{g}_{\rm ff}$ denotes the thermally-averaged Gaunt factor, which does the quantum mechanical correction to the classical electrodynamics. Its value varies between $1$ to $5$, depending on the electron energy. However, following Yarza et al. (2020), we take $\bar{g}_{\rm ff}=1.2$ throughout this paper. Note that the second term in Eq. (24) includes both the relativistic corrections and electron-electron bremsstrahlung emission. We also assume a two-temperature plasma in which the electron temperature is lower than that of the ions. Here, we briefly discuss the physical reasons behind this assumption. In an ionized plasma, electrons and ions attempt to maintain thermal equilibrium through random Coulomb collisions. However, this equilibrium breaks down, particularly near the horizon. In this region, the flow velocity becomes extremely high, and the accretion timescale becomes much shorter than the electron-ion collision timescale. As a result, the assumption of a single-temperature plasma becomes unrealistic, and the accretion flow tends to enter a two-temperature regime, especially in the inner regions of the disc. There are many studies in the literature corresponding to the two-temperature accretion flows Sarkar and Chattopadhyay (2018); Dihingia et al. (2018b, 2020a); Sarkar et al. (2020); Sarkar and Chattopadhyay (2022), where the electron and ion temperature profiles are obtained self-consistently by solving the radial momentum and energy equations for each species. However, in our work, we adopt a simplified scaling relation, $T_{e}=T/10$ in accordance with the works Mościbrodzka et al. (2009); Yarza et al. (2020). In those studies, the authors employed this scaling relation to successfully model the observed data of Sgr A*. However, readers may adopt some alternative scaling relations as discussed in works Chattopadhyay and Chakrabarti (2002); Dihingia et al. (2020b).
The emission coefficient of accretion flow, as described in Eq. (24), is measured locally. However, for an observer at spatial infinity, this radiation must be red-shifted due to both the gravitational potential of the black hole and the Doppler effect resulting from the rotation of the disc. Here, we do not consider the light bending effect to the photons for simplicity. Also, the disc inclination angle with respect to the line of sight of the observer is taken as $45^{\circ}$ Dihingia et al. (2020b); Sen et al. (2022); Patra et al. (2024a). Under these assumptions, the observed frequency $\nu_{o}$ is obtained in terms of the emitted frequency $\nu_{e}$ as Luminet (1979),
@@
$\begin{split}L_{\nu_{o}}&=2\int_{x_{0}}^{x_{\rm edge}}\int_{0}^{2\pi}\mathcal{E}_{\nu_{e}}^{\rm ff}Hxdxd\phi\\ &=\frac{64\pi e^{6}}{3m_{e}m_{p}^{2}c^{3}}\sqrt{\frac{2\pi}{3m_{e}k_{B}}}\bar{g}_{\rm ff}\int_{x_{0}}^{x_{\rm edge}}\int_{0}^{2\pi}\biggl[Hx\rho^{2}T_{e}^{-1/2}\\ &\times(1+4.4\times 10^{-10}T_{e})\exp\left(-\frac{(1+z)h\nu_{o}}{k_{B}T_{e}}\right)\biggr]dxd\phi\\ &\times{\text{erg}~\text{s}^{-1}~\text{Hz}^{-1}},\end{split}$ (27)
where $x_{0}$ ($x_{\rm H}=$ event horizon radius) and $x_{\rm edge}$ are the inner and outer edges of the CPD, respectively.
Bremsstrahlung radiation can only be observed when the emitting medium is optically thin; otherwise, photons will be destroyed through true absorption after undergoing multiple coherent scattering. The primary source of opacity is the Thomson scattering by the free electrons with opacity coefficient $\kappa_{s}=0.4~{\text{cm}^{2}~\text{gm}^{-1}}$ Maoz (2016). In a fully ionized medium, free-free absorption also significantly contributes to the photon opacity Maoz (2016). Using Eq. (24), we calculate the Rosseland mean opacity coefficient for the free-free absorption as Shapiro and Teukolsky (2008),
@@
Here, we consider the typical length scale of the medium is the half-thickness ($H$) of the disc. Therefore, the effective optical depth of the disc is obtained by considering the aforementioned opacity sources as Rybicki and Lightman (1991),
@@
where $\tau_{s}$ ($\kappa_{s}\rho H$) is the scattering optical depth and $\tau_{a}$ ($\kappa_{\rm R}^{\text{ff}}\rho H$) is the absorption optical depth.
@@ IV.1 Analysis of critical points
We already pointed out that the transonic accretion flow must pass through the critical points; therefore, we start our analysis by investigating how the behaviors of critical points depend on different global constants, such as specific angular momentum ($\lambda$), binary mass ratio ($q$), and binary separation ($z_{2}$). To do this, we first compute the flow temperature ($\Theta_{c}$) and radial velocity ($v_{c}$) at a critical point ($x_{c}$) for a given set of input parameters ($\lambda,q,z_{2}$) by solving the critical point conditions (23). Subsequently, using these results in Eq. (10), we calculate the specific energy ($E$) at $x_{c}$ for the same set of input parameters mentioned above. The obtained results are shown in Fig. 3, where the variation of $E$ as a function of $x_{c}$ is illustrated. In panel (a), we fix $(q,z_{2})=(1,500)$ and vary the angular momentum as $\lambda=3.3,3.35$, and $3.4$. In panel (b), we set $(\lambda,z_{2})=(3.5,500)$ and vary the mass ratio as $q=0.01,0.1$, and $1$. Finally, in panel (c), we choose $(\lambda,q)=(3.5,0.5)$ and vary the binary separation as $z_{2}=500,750$, and $1000$. The color codes corresponding to the different input parameters are indicated in each panel. In each curve, the saddle, nodal, and spiral-type critical points are represented by solid, dotted, and dashed lines, respectively. We observe that the flow energy remains constant in all cases when the critical points are located at large distances from the black hole. However, the flow energy changes significantly when the critical points are closer to the black hole. For each curve, the critical points are formed in a systematic order as saddle-nodal-spiral-nodal-saddle-spiral when the critical points move toward the outer edge of the disk. Moreover, we observe that the flow energy changes moderately with an increase in $\lambda$. However, it shows small variation with increases in $q$ and $z_{2}$. Interestingly, a broad range of energy values exists where the flow possesses two saddle-type critical points (as indicated by the horizontal dashed lines (magenta) in all the panels). It is important to note that such regions with multiple critical points are of particular interest in the study of shock-induced accretion solutions. For the shock scenarios Dihingia et al. (2018a); Dihingia et al. (2019); Patra et al. (2022); Sen et al. (2022); Mitra and Das (2024); Patra et al. (2024b), a global solution can pass through both the inner (i.e., formed near the black hole) and outer (i.e., formed far from the black hole) critical points, provided that the relativistic shock conditions are satisfied. A detailed analysis of the shock solutions is presented in Section IV.3.
@@ IV.2 Effect of binary parameters on the transonic accretion solutions
In this section, we aim to investigate the effect of the secondary black hole (SBH) on the accretion solutions in the CPD. To determine the transonic accretion solutions, we calculate the critical point $x_{c}$ and its corresponding flow variables $v_{c}$ and $\Theta_{c}$ for a given set of input parameters ($q,z_{2},\lambda,E$) by using Eqs. (23) and (10). Next, using the calculated parameters ($\Theta_{c},v_{c}$) at $x_{c}$ as the initial boundary conditions, we first integrate Eqs. (18) and (20) from $x_{c}$ to the disc inner edge $x_{0}$, and then from $x_{c}$ to the disc outer edge $x_{\text{edge}}$. Finally, by combining both segments of the solutions, we obtain the resulting global accretion solution corresponding to the same set of input parameters ($q,z_{2},\lambda,E$). In this work, the inner and outer edges of the accretion disc are taken as $x_{0}=0.05$ and $x_{\text{edge}}=350$, respectively. It is essential to mention that there is no qualitative difference in the characteristics of the accretion solutions if the inner edge position is chosen more closer to the horizon ($x_{\rm H}=0$) instead of $x_{0}=0.05$. In Fig. 4, we depict the typical accretion solutions (i.e., variation of the Mach number ($M=v/C_{s}$) as a function of the radial coordinate ($x$)) for different values of $q$ and $z_{2}$. Here, we choose the input flow parameters as $\lambda=3.5$ and $E=1.01$. The left panels represent the accretion solutions corresponding to the mass ratio $q=0.01,0.1$, and $1$ with binary separation $z_{2}=500$. Similarly, the right panels denote the accretion solutions for $z_{2}=500,750$, and $1000$ with $q=0.5$. In all cases, the solutions that pass through the outer critical points ($x_{\text{out}}$) are globally extended from $x_{\text{edge}}$ to $x_{0}$. In contrast, the solutions that pass through the inner critical points ($x_{\text{in}}$) are truncated at certain radii ($x_{t}$) in between $x_{\text{in}}$ and $x_{\text{out}}$. Accretion solutions of this kind are usually referred as A-type solutions Dihingia et al. (2018a, 2020b); Sen et al. (2022); Patra et al. (2022, 2024a, 2024b). Interestingly, for A-type solution topology, the solution that passes through $x_{\text{out}}$ may connect to the solution that passes through $x_{\text{in}}$ via a standing shock transition, as the inner solution branch has a higher entropy content than the outer one Becker and Kazanas (2001). The shock-induced accretion solutions will be examined latter in Section IV.3. It is evident from the figure that the nature of the accretion solutions does not change with an increase in either of the binary parameters $q$ and $z_{2}$. The positions of the critical points associated with these accretion solutions are summarized in Table 1. This table shows that $x_{\text{in}}$ increases, while $x_{\text{out}}$ decreases as $q$ increases at a given $z_{2}$. On the other hand, when $z_{2}$ increases at a fixed $q$, $x_{\text{in}}$ and $x_{\text{out}}$ move oppositely compared to the previous case.